The area of a trapezoid is 110. If one base measures 9 and the height of the trapezoid is 10, find the length of the missing base.
Answer by stanbon(75887)

The area of a trapezoid is 110. If one base measures 9 and the height of the trapezoid is 10, find the length of the missing base.
-----------
Area = (1/2)h(b1+b2)
----
110 = (10/2)(9 + b)
----
220 = 10(9+b)
22 = 9+b
b = 13
===========
cheers,
Stan H.